OpenStudy (anonymous):

determine the limit of the integral [(3x-1)/(2n+3)]^(2x) from 1 to infinity, using the root test.

OpenStudy (anonymous):

is that series? integral ? what ? i guess \[\large \lim_{n \rightarrow \infty} (\frac{3n-1}{2n+3})^{2n}\]

OpenStudy (anonymous):

the sum from n=1 to infinity (or the integral both ways the limit is the same)

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

but yes we are talking about finding the sum of a series using the root test

OpenStudy (anonymous):

well Using root test \[a_{n}= (\frac{3n-1}{2n+3})^{2n}\] \[\large \lim_{n \rightarrow \infty} (a_{n})^{\frac{1}{n}}=\lim_{n \rightarrow \infty} [(\frac{3n-1}{2n+3})^{2n}]^{\frac{1}{n}}=\lim_{n \rightarrow \infty} (\frac{3n-1}{2n+3})^{2}=\frac{9}{4} > 1\]

OpenStudy (anonymous):

since we know that \[\lim_{n \rightarrow \infty} \frac{3n-1}{2n+3}=\frac{3}{2}\] so the series \[\large \sum_{n=1}^{\infty} (\frac{3n-1}{2n+3})^{2n}\] diverges

OpenStudy (anonymous):

ok thanks! i just forgot to square the answer i was putting in 3/2

OpenStudy (anonymous):

very well :)